Find the value of x in the inequality Ix-7I is less then or equal to 2.
NOTE-II are absolute value signs.
Thanks,
Pat
: Find the value of x in the inequality Ix-7I is less then or equal to 2.
NOTE-II are absolute value signs.
Thanks,
Pat

Answer by Earlsdon(3717)

Find the value of x:
abs(x-7)<= 2 Remove the absolute-value bars to get the two inequalitites:
x-7 <= 2 and x-7 >= -2 Now solve each of these to get:
x<=9 and x>=5